# The entries of the Sinkhorn limit of an $m\times n$ matrix

### v1

@@ 4. Rectangular matrices and combinatorial structure
Let $m=2$ and $S=\begin{smallmatrix}\{\}&\{2\}&\{2\}&\cdots&\{2\}\\ \{\}&\{2\}&\{3\}&\cdots&\{k\}\end{smallmatrix}$. Then
$\operatorname{adj}_{S}(2,n)=\begin{bmatrix}-2n&2&2&\cdots&2\\ -n&2-n&-2&\cdots&-2\\ -n&-2&2-n&\cdots&-2\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ -n&-2&-2&\cdots&2-n\end{bmatrix}.$
If $k\in\{1,2\}$, then $2^{\lvert{S}\rvert}c_{S}(2,n)=\det\operatorname{adj}_{S}(2,n)$, so these signs are correct. If $k\geq 3$, they are not, but negating each $-2$ entry (that is, the off-diagonal entries that aren’t in the first row or first column) gives a matrix whose determinant is $2^{\lvert{S}\rvert}c_{S}(2,n)$.
Example 32.
Let $m=2$ and $T=\begin{smallmatrix}\{2\}&\{2\}&\{2\}&\cdots&\{2\}\\ \{2\}&\{3\}&\{4\}&\cdots&\{k+1\}\end{smallmatrix}$. Then
$\operatorname{adj}_{T}(2,n)=\begin{bmatrix}2-n&-2&\cdots&-2\\ -2&2-n&\cdots&-2\\ \vdots&\vdots&\ddots&\vdots\\ -2&-2&\cdots&2-n\end{bmatrix}.$
If $k\in\{0,1,2\}$, then these signs are correct. If $k\geq 3$, then (again) negating the off-diagonal entries that aren’t in the first row or first column gives a matrix whose determinant is $2^{\lvert{S}\rvert}c_{T}(2,n)$.
The determinants of the two previous sign-corrected matrices evaluate as follows.
@@
(We have defined $c_{S_{0}}(2,n)=0$ and $c_{T_{n}}(2,n)=0$ despite $S_{0}$ and $T_{n}$ being undefined; this allows us to write the following sum simply.) For every positive $2\times n$ matrix $A$, the top left entry $x$ of $\operatorname{Sink}(A)$ satisfies
@@ 5. Open questions
The main open question is to identify the equivalence class of correct sign alterations $\sigma(S)$ in Conjecture 2 for each $S$, ideally in the form of a canonical representative. Since $\operatorname{adj}_{S}(m,n)$ resembles an adjacency matrix, it seems likely that $\det\operatorname{adj}_{S,\sigma(S)}(m,n)$ has a combinatorial interpretation, and this may suggest the correct signs. A natural candidate is a signed graph, which is a graph along with a function assigning $-1$ or $1$ to each edge. The switching class of a signed graph is the set of all signed graphs that can be obtained by choosing a subset $V$ of its vertices and negating the signs of all edges incident to a vertex in $V$ (with multiplicity if an edge is incident to multiple vertices in $V$[24]. The equivalence class of correct sign alterations for $S$ corresponds to a switching class on the graph $G_{S}$ defined in Section 4. We haven’t been able to use this to identify the correct signs, however.
Second, how can we prove Conjecture 2? In the absence of a combinatorial proof, one could interpolate polynomial equations representing the entries of diagonal matrices $R$ and $C$, as we did for the entries of $\operatorname{Sink}(A)$, and hope to prove that all three matrices are correct by checking that they satisfy $\operatorname{Sink}(A)=RAC$. This is how Nathanson established the Sinkhorn limit of a $2\times 2$ matrix [16, Theorem 3]. However, this works for $2\times 2$ matrices because we have explicit expressions for the entries of $\operatorname{Sink}(A)$, rather than specifications as roots. For $3\times 3$ matrices, the diagonal entries of $R$ and $C$ are roots of degree-$6$ polynomials, and generically the product of two such roots has degree $6^{2}$; this is too big, since the entries of $\operatorname{Sink}(A)$ have degree $6$. Therefore we would need to find a single degree-$6$ field extension that contains all entries of $\operatorname{Sink}(A)$, $R$, and $C$ so that we can perform arithmetic on them symbolically. What is this extension?
Third, as we saw in Section 2, the lack of a unique representation of the coefficient of $x^{3}$ for $3\times 3$ matrices is due to a relation among $12$ monomials $M(S)$ with $\lvert{S}\rvert=3$. What is the combinatorial structure of such relations? Similarly, is there structure in relations among $\Sigma(S)$? For example, the coefficient of $x^{5}$ for $3\times 4$ matrices has a $1$-dimensional family of representations due to the relation
@@
Fourth, how does Corollary 8 generalize to $m\times n$ matrices with linear dependencies among their rows or columns? The coefficients in Corollary 8 have unique representations as linear combinations of class sums $\Sigma(S)$ where $S\subseteq D(2,3)$, namely
@@
Finally, there is a further generalization of Conjecture 2 whose form is not known. Decades before Sinkhorn’s paper, the iterative scaling process was introduced by Kruithof [14] in the context of predicting telephone traffic. In this application, rather than scaling to obtain row and column sums of $1$, each row and column has a potentially different target sum. Sinkhorn [20] showed that the limit exists. We call this limit the Kruithof limit.
@@
with target row sums $V=\begin{bmatrix}6000&4000&2500&1000\end{bmatrix}^{\textnormal{{T}}}$ and target column sums $W=\begin{bmatrix}6225&4000&2340&935\end{bmatrix}$. Let $x$ be the top left entry of the Kruithof limit. Numerically, $x\approx 3246.38700234$. A Gröbner basis computation gives an equation
@@
The previous example suggests that the degrees of entries of Kruithof limits are the same as those of Sinkhorn limits of the same size.
@@
Let $m\geq 1$ and $n\geq 1$. Let $A$ be a positive $m\times n$ matrix, let $V$ be a positive $m\times 1$ matrix, and let $W$ be a positive $1\times n$ matrix such that the sum of the entries of $V$ equals the sum of the entries of $W$. The top left entry $x$ of the Kruithof limit of $A$ with target row sums $V$ and target column sums $W$ is algebraic over the field generated by the entries of $A$, $V$, and $W$, with degree at most $\binom{m+n-2}{m-1}$.
Since the Kruithof limit specializes to the Sinkhorn limit when $V=\begin{bmatrix}1&1&\cdots&1\end{bmatrix}^{\textnormal{{T}}}$ and $W=\begin{bmatrix}\frac{m}{n}&\frac{m}{n}&\cdots&\frac{m}{n}\end{bmatrix}$, we suspect that the entries in the determinant formulas in Conjecture 27 generalize somehow to involve entries of $V$ and $W$, and this should give a generalization of Conjecture 2 to Kruithof limits. Moreover, Gröbner basis computations suggest that the surprising property we mentioned in Section 3 regarding the coefficients for square matrices satisfying
$\sum_{S\subseteq D(n,n)}c_{S}(n,n)x^{\lvert{S}\rvert}=(x-1)^{\binom{2n-2}{n-1}}$
generalizes as follows.
@@
With the notation of Conjecture 35, let $V=\begin{bmatrix}r_{1}&r_{2}&\cdots&r_{m}\end{bmatrix}^{\textnormal{{T}}}$ and $W=\begin{bmatrix}c_{1}&c_{2}&\cdots&c_{n}\end{bmatrix}$. If the general equation satisfied by the top left entry $x$ of the Kruithof limit of $A$ with target row sums $V$ and target column sums $W$ is
@@
then

### v2

@@ 4. Rectangular matrices and combinatorial structure
The subsets $S$ and $T$ in Example 29 each consist of $5$ type-2 links. For $S$, we have $C_{1}=\{2,3,4\}$ to $C_{2}=\{2,3,5\}$. The element $4\in C_{1}$ plays the role of $5\in C_{2}$, so we consider the permutation mapping $(2,3,s)$ to $(2,3,s)$. This is the identity permutation, which has sign $1$, and indeed the $(1,2)$ and $(2,1)$ entries in the first matrix in Example 29 have sign $1$. Analogously, the signs in the $5$-cycle
$\{2,3,4\}\mapsto\{2,3,5\}\mapsto\{2,5,7\}\mapsto\{2,6,7\}\mapsto\{2,4,6\}\mapsto\{2,3,4\}$
are $1,-1,1,-1,-1$. These agree with the signs of the remaining off-diagonal entries. On the other hand, for the subset $T$, the signs in the $5$-cycle
$\{2,3,4\}\mapsto\{2,3,5\}\mapsto\{3,5,6\}\mapsto\{4,5,6\}\mapsto\{2,4,6\}\mapsto\{2,3,4\}$
are $1,1,1,-1,-1$. These agree with the signs in the second matrix in Example 29.
The sign of each permutation can be encoded more simply by the position where the differing index $s$ appears in each ordered list. This formulation gives the definition of $\operatorname{adj}_{S}(m,n)$ in Section 1 and results in Conjecture 2.
We conclude this section by specializing Conjecture 2 to $2\times n$ matrices. For each exponent $k$, there are at most $2$ equivalence classes of subsets — an equivalence class containing subsets $S$ such that ({},{}$(\{\},\{\})\in S$ and another containing the rest. Evaluating the two determinant formulas gives the following.
@@
(We have defined $c_{S_{0}}(2,n)=0$ and $c_{T_{n}}(2,n)=0$ despite $S_{0}$ and $T_{n}$ being undefined; this allows us to write the following sum simply.) For every positive $2\times n$ matrix $A$, the top-left entry $x$ of $\operatorname{Sink}(A)$ satisfies
@@ 5. Open questions
We mention several open questions. The main question is how to prove Conjecture 2. Until we understand why the two types of links arise, a simple proof seems unlikely. Alternatively, toward an algebraic proof, one could interpolate polynomial equations satisfied by the entries of diagonal matrices $R$ and $C$, as we did for the entries of $\operatorname{Sink}(A)$, and hope to prove that all three matrices are correct by checking that they satisfy $\operatorname{Sink}(A)=RAC$. This is how Nathanson established the Sinkhorn limit of a $2\times 2$ matrix [16, Theorem 3]. However, this works for $2\times 2$ matrices because we have explicit expressions for the entries of $\operatorname{Sink}(A)$, rather than specifications as roots. For $3\times 3$ matrices, the diagonal entries of $R$ and $C$ are roots of degree-$6$ polynomials, and generically the product of two such roots has degree $6^{2}$; this is too big, since the entries of $\operatorname{Sink}(A)$ have degree $6$. Therefore we would need to find a single degree-$6$ field extension that contains all entries of $\operatorname{Sink}(A)$, $R$, and $C$ so that we can perform arithmetic on them symbolically. What is this extension?
Second, as we saw in Section 2, the lack of a unique representation of the coefficient of $x^{3}$ for $3\times 3$ matrices is due to a relation among $12$ monomials $M(S)$ with $\lvert{S}\rvert=3$. What is the combinatorial structure of such relations? Similarly, is there structure in relations among $\Sigma(S)$? For example, the coefficient of $x^{5}$ for $3\times 4$ matrices has a $1$-dimensional family of representations due to the relation
@@
Third, how does Corollary 8 generalize to $m\times n$ matrices with linear dependencies among their rows or columns? The coefficients in Corollary 8 have unique representations as linear combinations of class sums $\Sigma(S)$ where $S\subseteq D(2,3)$, namely
@@
Fourth, why does Conjecture 2 imply Conjectures 10 and 11?
Finally, there is a generalization of Conjecture 2 for which the determinant formula is not known. Decades before Sinkhorn’s paper, the iterative scaling process was introduced by Kruithof [14] in the context of predicting telephone traffic. In this application, rather than scaling to obtain row and column sums of $1$, each row and column has a potentially different target sum. Sinkhorn [21] showed that the limit exists. We call this limit the Kruithof limit.
@@
with target row sums $V=\begin{bmatrix}6000&4000&2500&1000\end{bmatrix}^{\textnormal{{T}}}$ and target column sums $W=\begin{bmatrix}6225&4000&2340&935\end{bmatrix}$. Let $x$ be the top-left entry of the Kruithof limit. Numerically, $x\approx 3246.38700234$. A Gröbner basis computation gives the equation
@@
The previous example suggests that the degrees of entries of Kruithof limits are the same as those for Sinkhorn limits with the same dimensions.
@@
Let $m\geq 1$ and $n\geq 1$. Let $A$ be a positive $m\times n$ matrix, let $V$ be a positive $m\times 1$ matrix, and let $W$ be a positive $1\times n$ matrix such that the sum of the entries of $V$ equals the sum of the entries of $W$. The top-left entry $x$ of the Kruithof limit of $A$ with target row sums $V$ and target column sums $W$ is algebraic over the field generated by the entries of $A$, $V$, and $W$, with degree at most $\binom{m+n-2}{m-1}$.
Since the Kruithof limit specializes to the Sinkhorn limit when $V=\begin{bmatrix}1&1&\cdots&1\end{bmatrix}^{\textnormal{{T}}}$ and $W=\begin{bmatrix}\frac{m}{n}&\frac{m}{n}&\cdots&\frac{m}{n}\end{bmatrix}$, we suspect that the entries of $\operatorname{adj}_{S}(m,n)$ generalize in some way to involve entries of $V$ and $W$, and this should give a generalization of Conjecture 2 to Kruithof limits. What are the generalized entries?
We mention that Conjecture 11 seems to generalize to Kruithof limits. For example, the top-left entry $x$ of the Kruithof limit of a general positive $2\times 3$ matrix $A$ with target row sums $V=\begin{bmatrix}r_{1}&r_{2}\end{bmatrix}^{\textnormal{{T}}}$ and target column sums $W=\begin{bmatrix}c_{1}&c_{2}&r_{1}+r_{2}-c_{1}-c_{2}\end{bmatrix}$ satisfies $f_{3}x^{3}+f_{2}x^{2}+f_{1}x+f_{0}=0$, where
$f_{3}$ $=M\!\left(\!\begin{smallmatrix}\{\}&\{2\}&\{2\}\\ \{\}&\{2\}&\{3\}\end{smallmatrix}\!\right)$
$f_{2}$ $=(c_{2}-r_{1}-r_{2})M\!\left(\!\begin{smallmatrix}\{\}&\{2\}\\ \{\}&\{2\}\end{smallmatrix}\!\right)-(c_{1}+c_{2})M\!\left(\!\begin{smallmatrix}\{\}&\{2\}\\ \{\}&\{3\}\end{smallmatrix}\!\right)+(r_{2}-c_{1})M\!\left(\!\begin{smallmatrix}\{2\}&\{2\}\\ \{2\}&\{3\}\end{smallmatrix}\!\right)$
$f_{1}$ $=c_{1}(r_{1}+r_{2})M\!\left(\!\begin{smallmatrix}\{\}\\ \{\}\end{smallmatrix}\!\right)+c_{1}(r_{1}-c_{2})M\!\left(\!\begin{smallmatrix}\{2\}\\ \{2\}\end{smallmatrix}\!\right)+c_{1}(c_{1}+c_{2}-r_{2})M\!\left(\!\begin{smallmatrix}\{2\}\\ \{3\}\end{smallmatrix}\!\right)$
$f_{0}$ $=-r_{1}c_{1}^{2}M\!\left(\!\begin{smallmatrix}\vphantom{\{}\\ \vphantom{\{}\end{smallmatrix}\!\right).$
The coefficients of the monomials $M(S)$ in the expressions for $f_{k}$ satisfy
$x^{3}+\big{(}(c_{2}-r_{1}-r_{2})-(c_{1}+c_{2})+(r_{2}-c_{1})\big{)}x^{2}\\ +\big{(}c_{1}(r_{1}+r_{2})+c_{1}(r_{1}-c_{2})+c_{1}(c_{1}+c_{2}-r_{2})\big{)}x-r_{1}c_{1}^{2}\\ =(x-r_{1})(x-c_{1})^{2}.$
Additional Gröbner basis computations suggest the following.
@@
With the notation of Conjecture 33, write $V=\begin{bmatrix}r_{1}&r_{2}&\cdots&r_{m}\end{bmatrix}^{\textnormal{{T}}}$ and $W=\begin{bmatrix}c_{1}&c_{2}&\cdots&c_{n}\end{bmatrix}$ where $r_{1}+r_{2}+\dots+r_{m}=c_{1}+c_{2}+\dots+c_{n}$. If the general equation satisfied by the top-left entry $x$ of the Kruithof limit of $A$ with target row sums $V$ and target column sums $W$ is
@@
where the equation is scaled so that $c_{D(m,n)}(V,W)=1$, then